A traffic survey on a busy highway reveals that one of every four vehicles is a truck. This survey also established that one-eighth of all automobiles are unsafe to drive and one-twentieth of all trucks are unsafe to drive.
a. What i s the probability that the next vehicle to pass a given point i s an unsafe truck?
b. What is the probability that the next vehicle will be a truck, given that it is unsafe?
c. What is the probability that the next vehicle that passes a given point will be a truck, given that the previous vehicle was an automobile?
1
Expert's answer
2020-12-24T12:51:05-0500


P(T) = probability of truck

P(U) = probability of unsafe vehicle

"P(T) = \frac{1}{4} \\

P(R) = \frac{3}{4} \\

P(U) = \frac{1}{8} \\

P(U/T) = \frac{1}{20} \\

P(U) = P(T)P(U/T) + P(R)P(U/R) \\

\frac{1}{8} = \frac{1}{4} \times \frac{1}{20} + \frac{3}{4}P(U/R) \\

P(U/R) = \frac{3}{20}"

a.

"P(U \cap T) = P(T)P(U/T) \\

= \frac{1}{4} \times \frac{1}{20} \\

= \frac{1}{80} \\

= 0.0125"

b.

"P(T/U) = \frac{P(U \cap T)}{P(U)} \\

= \frac{1}{80} \times \frac{8}{1} \\

= \frac{1}{10} \\

= 0.1"

c.

"P(T/A) = \frac{P(T \cap A)}{P(A)} \\

= \frac{P(T)}{P(A)} \\

= \frac{\frac{1}{4}}{1} \\

= \frac{1}{4} \\

= 0.25"